# Equal contributions

### v2

@@ Abstract
Run-by-run variability in parallel programs caused by floating-point non-associativity (FPNA) has been known to significantly affect reproducibility in iterative algorithms, due to accumulating errors. Non-reproducibility negatively affects efficiency and effectiveness of correctness testing for stochastic programs. Recently, the sensitivity of deep learning (DL) training and inference pipelines to FPNA have been found to be extreme, and can prevent certification for commercial applications, accurate assessment of robustness and sensitivity, and bug detection. New approaches in scientific computing applications have coupled DL models with high-performance computing (HPC) simulations, leading to an aggravation of debugging and testing challenges. Here we perform an investigation of the statistical properties of FPNA within modern parallel programming models, analyze performance and productivity impacts of replacing atomic operations with deterministic alternatives on GPUs, and examine the recently-added deterministic options within the PyTorch framework within the context of GPU deployment, uncovering and quantifying the impacts of input parameters triggering run-by-run variability and reporting on the reliability and completeness of the documentation. Finally, we evaluate the strategy of exploiting automatic determinism provided by deterministic hardware, using the Groq $\text{LPU}^{\text{TM}}$ accelerator for inference portions of the DL pipeline. We demonstrate the benefits that this strategy can provide within reproducibility and correctness efforts.
@@ I Introduction
Run-by-run variability of a program, despite identical inputs and software stack, is usually assumed to be negligible, even in heterogeneous parallel programming [1]. However, variability caused by FPNA coupled with asynchronous parallel operations such as reductions can be substantial [2], especially for massively parallel programs using iterative stochastic routines, such as those implementing optimization algorithms like conjugate gradient [3], and can mask errors within threshold-based correctness testing schemes [4], which are often used in scientific computing programs such as molecular simulation [5, 6, 7], making debugging difficult. Deep neural network (DNN) training also involves iterative, stochastic algorithms coupled with non-linear activation functions, and this combination has been found to cause extreme sensitivity to bit-level numerical changes, such as those caused by FPNA [8, 9, 10]. DNN inference can also suffer from such sensitivity, due to the influence of non-linear activation functions, although the effects may be reduced due to the absence of the compounding errors caused by the iterative training schemes. Recently, a number of studies have shown that this run-by-run variability in the full DNN training and inference pipeline can lead to unacceptably large differences in the predictions produced by a model, and have thwarted efforts to release reproducible commercial applications in safety-critical sectors such as medical diagnostics [11] and autonomous driving [8]. In addition, recent reports have found that all of the major DL software frameworks such as TensorFlow and PyTorch contain hundreds of bugs originating from all software levels, including environment misuse, incorrect assignment, and concurrency issues such as data races [12]; bugs in these frameworks can be disastrous [13], especially when they are silent [14], and high run-by-run variability can make it extremely difficult to detect bugs in these deep, multi-language, multi-level parallel stacks.
@@
Within HPC scientific computing, the incorporation of deep learning into traditional approaches for simulation has become increasingly popular [15, 16, 17]. Within molecular simulation for example, DNN models for interatomic potentials have promised quantum mechanical accuracy at the cost of simpler, empirical models, which translates to speedups of several orders of magnitude; this approach received the ACM Gordon Bell Award in 2020 [18]. Our preliminary tests of reproducibility when using identical inputs for the training and inference pipeline for these types of models revealed a level of non-reproducibility for prediction of forces on atoms that would be unacceptable for the traditional quantum mechanical HPC programs [19]. Besides asynchronous parallel reductions within the deep learning framework itself, external parallel kernels programmed by the simulation developers that feed into training pipelines [18] can also introduce run-by-run reproducibility.
@@
II Experimental Methods
In the following we describe the hardware and systems software used to perform our numerical analyses and the mathematical definitions developed to quantify output variability caused by non-determinism due to FPNA in parallel kernels, studied as isolated test cases and within the PyTorch deep learning framework high-level operations.
II-A Hardware description
Tests on GH200 are run on the Alps supercomputing system located at CSCS running SLE 15 (enterprise). Each node has 4 GH200 with 72 ARM cores, 128 GB LPDDR5, and one H100 GPU with 96 GB HBM3 memory each. Tests for the V100 are run on the Summit supercomputer at the Oak Ridge Leadership Computing Facility (OLCF), running Redhat OS 8. Summit is an IBM system; each IBM Power System AC922 node has two Power9 CPUs with 512 GB of memory and 6 V100 NVidia GPU with 16GB of HBM2 memory. Tests on Mi250X AMD GPU are obtained on the Frontier supercomputer at OLCF, running SLE 15 (enterprise). Frontier is an HPE Cray EX supercomputer; each Frontier compute node has a 64-core AMD“Optimized 3rd Gen EPYC” CPU with 512 GB of DDR4 memory and 4 AMD MI250X GPUs, each with 2 Graphics Compute Dies (GCDs) for a total of 8 GCDs per node. Tests on the NVidia H100 are performed on a Groq host node3 in Groq ZankerLab running Ubuntu LTS 22.04.06. This machine has 2 H100’s with 40GB HBM3 memory each and has an AMD EPYC 7302 16-Core Processor host CPU with 2 sockets, 16 cores per socket and 2 threads per core. Tests on the LPU accelerator are run on the $\text{GroqNode}^{\text{TM}}$ server r01-gn-01 in Groq ZankerLab with $\text{GroqWare}^{\text{TM}}$ SDK 0.11 and Ubuntu LTS 22.04. A GroqNode has 8 Groq LPUs connected with a fully connected 88 RealScale™ chip-to-chip connectors and 2 AMD EPYC 7313 processors (3GHz, 16C/32T, 155W TDP each).
II-B Metrics for measuring non-determinism in operations
We defined metrics $V$ for quantifying the run-by-run variability in output values for operations with scalar and multi-dimensional array inputs/outputs, following a similar approach used in error analysis [20]. These are explained below. In all cases, $V$ is a normalized quantity, with $V=0$ if two implementations are bitwise identical, nonzero otherwise, and with $V=1$ being the highest possible variability.
II-B1 Scalar-valued outputs
We use $V_{s}(f)=1-f_{\textsc{nd}}/f_{\textsc{d}}$ to quantify the bitwise non-determinism between two implementations of some function $f$, where the subscripts nd and d label the non-deterministic and deterministic implementations respectively.
II-B2 Array outputs
To quantify the bitwise non-determinism between two implementations of a function producing a multidimensional array output, we define two different metrics. Let two implementations of a function $f$ be $f_{1}$ and $f_{2}$, which produce as outputs the arrays $\mathbf{A}$ and $\mathbf{B}$, respectively, each with dimensions $d_{1},d_{2},\ldots,d_{k}$ and $D$ total elements. The first metric is the elementwise relative mean absolute variation,
@@ II-B2 Array outputs
The second metric, the count variability, measures how many elements in the multidimensional array are different,
@@
III Impact of atomic operations on parallel reductions: the recursive sum example
Floating point operations, being non-associative, introduce variability in outputs when computed in different orderings. To illustrate the magnitude of this variability, we use the recursive sum operation, $S$. Consider the following sum $S_{\textsc{d}}=\sum_{i}^{n}x_{i}$ that we evaluate with a deterministic algorithm (the summation order is always the same). The simplest implementation adds the floating-point numbers (FP) $x_{i}$ in the order they are stored in serial (thus the “recursive” nomenclature). When parallelized with asynchronous operations and executed in an unspecified order, this is equivalent to applying a random permutation $P$ to the series $x_{i}$ before computing the serial sum, $S_{\textsc{nd}}=\sum_{i}^{n}x_{P(i)}$.
@@ III Impact of atomic operations on parallel reductions: the recursive sum example
We generated lists $x_{i}$ of double precision floating point numbers (FP64) numbers of various lengths using Python, with $x_{i}\in\mathcal{N}(0,1)$ the normal distribution of zero mean and $\sigma=1$, and computed $S_{\textsc{d}}$ and $S_{\textsc{nd}}$ before and after applying a random permutation $P$ to $x_{i}$, repeating ten times. As shown in Tab. I, the variability introduced by the permutation can be larger than the tolerance of some of the correctness tests included in the computational quantum mechanics CP2K[5], as some of the tests have tolerances as low as $10^{-14}$. This illustrates the problems that FPNA can introduce in correctness testing schemes for computational tasks with stringent requirements.
III-A Different parallel sum implementations on GPU
A difficulty for asynchronous programming on GPUs is the absence of a global synchronization barrier at the kernel level; to avoid races one can use (i) atomic instructions, (ii) the implicit global synchronization happening when multiple kernels are added to the same stream (iii) one GPU thread to calculate the sum, (iv) the __threadfence instruction.
1__global__ void reduce_atomic_only(double *sdata, int size, double *res) {
@@ III-A Different parallel sum implementations on GPU
5}
6
7__device__ void block_reduce(double *smem, double *sdata, int tid, int size) {
@@
17}
18
19__global__ void reduce_pairwise(double *sdata, int size, double *res) {
20 int tid = threadIdx.x + blockIdx.x * blodkDim.x;
21 extern __shared__ double smem[];
22
23 block_reduce(smem, sdata, tid, size);
@@
26
27 // to avoid the second stage, we can replace the line above with
28 // if (threadIdx.x == 0) atomicAdd(res, smem[0]);
29}
Listing 1: Implementation of variants of parallel sums using the
atomicAdd instruction, and the pairwise reduction.
Option (iii) is deterministic but slow as it is performed in serial. Listing 1 shows the options for implementing a parallel sum on GPU. Using a single atomicAdd instruction, as illustrated by the function reduce_atomicadd_only of Listing 1, is an easy option from the programming perspective, and therefore is frequently adopted by developers.
This atomicAdd-only (AO) method is sequential even though it runs on GPU, and yet the instruction order is runtime dependent, making it non-deterministic. In the pairwise algorithm [20, 21], the elements $x_{i}$ are added in pairs $t_{i}=x_{i}+x_{i+n/2}$, $i=1,n/2$ an operation done in parallel on GPU. This process is repeated recursively $log_{2}(n)$ times on $t_{i}$ to get the sum. To avoid the synchronization barrier needed between each stage, we distribute data over each thread block [21] (see listing 2), and use the __syncthreads instruction for thread synchronization within the thread block. Shared memory is also used to improve data locality and performance.
Three different methods are tested for accumulating all partial results. The simple-pass-with atomicAdd (SPA), uses an atomicAdd instruction to accumulate all partial sums (see the comment in Listing 1). This is again a more simple solution from the programmatic perspective, but the implementation is not deterministic. The two-passes-with-final-reduction-on-CPU (TPRC) method uses the property that two kernels launched on the same stream will execute sequentially following the submission order, introducing a synchronization barrier between them, or between a kernel and a data transfer between GPU and CPU. We choose the data transfer, and compute the final sum on CPU using the sequential recursive method. TPRC is deterministic, but more sensitive to compiler optimizations because of vectorization.
1__device__ int retirementCount = 0;
2__global__ void reduce_gpu_full(double *sdata, int size, double *res) {
3 int tid = threadIdx.x + blockIdx.x * blodkDim.x;
4 extern __shared__ double smem[];
5
6 block_reduce(smem, sdata, tid, size);
7
8 if (threadIdx.x == 0) res[blockIdx.x] = smem[0];
@@
22}
Listing 2: Illustration of the full block reduction on a GPU. This
implementation is deterministic. We assume that each thread block
has 64 threads.
The last approach, given by listing 2, uses an integer counter to keep track of the completed thread blocks. The sum function from the CUB/HIPCUB library (CU) uses a similar technique. Each thread block increments this variable atomically before exiting. The last block incrementing the counter is responsible for the remaining reduction, and can be performed in two ways. The single-pass-with-tree-reduction (SPS) method uses the block reduction algorithm while listing 2 uses the recursive sum; we call this recursive variant the single-pass-with-final-recursive-sum-on-GPU (SPRC) method. SPS and SPRC are both deterministic by construction. SPRC and SPS are only possible if we use the __threadfence instruction to avoid data race conditions. This instruction ensures that all writes issued by the calling thread are finished before the calling thread runs the next instruction, and notifies all the other running threads that the memory operation is over. The instruction is not a global synchronization barrier.
III-B Statistical properties of the nondeterministic parallel sum on GPUs
It is often assumed that the variability due to FPNA can be described as Gaussian noise, but we could not find evidence supporting this in the literature. We therefore performed numerical experiments to estimate the probability density function (PDF) of $V_{s}(S)$. We generated a set of 100 arrays of one million FP64 elements each taken from the uniform distribution $U(1,10)$, applied SPA 10000 times to each array, and evaluated $V_{s}$, using SPS for the deterministic summation, and returning one million values for $V_{s}$. $S_{\textrm{d}}$ thus refers here to SPS while $S_{\textrm{nd}}$ refers to SPA. We then repeated the experiment, replacing the uniform distribution with a normal distribution of zero mean and standard deviation 1.
Figure 1: Probability density of $V_{s}$ for sums of 1 million FP64 numbers sampled from the normal and uniform distributions on the V100 The kernel parameters are fixed to $N_{t}=64$ and $N_{b}=7813$ for both SPS and SPA.
The resulting PDF s are shown in Fig.1 for both distributions for the V100. The PDF of the Mi250X and GH200 included in the project GitHub [22] are similar to the V100. The repository also shows the results for the uniform distribution for the Mi250X and V100 GPU on the same axis: the mean and standard deviation of $V_{s}$ are different between the two GPU types. We also calculated the dependence $\text{Max}\left|V_{s}\right|$ as a function of the array size $n$ for the same sequences used for Fig. 1. The data can be fit with a power law of the form $\beta n^{\alpha}$. $\text{Max}\left|V_{s}\right|$ is proportional to $\sqrt{n}$ when $x_{i}\in U(0,10)$. The exponent is larger for $x_{i}\in\mathcal{N}(0,1)$, showing that the range of the numbers also plays a role. Using a Kullback–Leibler divergence criterion (KL) analysis, we find that all PDF s for SPA converge towards a normal distribution whose mean and standard deviation depend on $x_{i}$ and the GPU family. This result supports the assumption that the variability due to FPNA for a sum with SPA may be considered Gaussian noise.
We next repeated the experiments replacing SPA with AO, using a sample size of 500000 sums, with results shown in Fig. 2 for the V100. The distribution is a multimodal distribution instead of a normal distribution, showing that the assumption of Gaussian noise is invalid in general.
Figure 2: PDF of the scalar variability $V_{s}$ when AO is used for the nondeterministic implementation.
III-C Performance comparisons
To illustrate the performance impact of these different strategies, we measured the time needed to compute 100 sums of 4194304 FP64 elements taken from the uniform distribution, using SPS, TPRC, CU, AO and SPA. The only nondeterministic algorithms in this tests are AO and SPA. We compute the negative quantity $P_{s}=100(1-\text{min}(t_{i})/t_{i})$ where $\text{min}(t_{i})$ is the minimum of all timings $t_{i}$, to compare the performance of all implementations to the fastest one. The other parameters controlling the algorithm are the thread block size and the number of thread blocks. Details on the choice of these settings are provided in the GitHub [22].
The main results are summarized on Tab. II. We find that AO is 4 orders of magnitude slower than the fastest implementations. The fastest sum implementation depends on the GPU version. The non-deterministic SPA seems to be favored on NVIDIA GPU, followed by SPS and CU. The difference in timings between SPA and SPS is less than 0.2% for the V100 GPU but can reach up to 7.8% on the GH200. CU suffers a 4.5% penalty on GH200. Most implementations on V100 are within 0.5% to 1% of each other. The performance penalty on GH200 is more spread than on V100. TPRC is the fastest implementation on the Mi250X GPU followed by CU.
TABLE II: Timings and performance penalty of various implementations of the summation algorithm on V100, GH200 and Mi250X GPU for 100 sums of 4194304 FP64 and different kernel parameter values. Timings are averaged over 10 consecutive runs.
GPU implementation ($N_{t}\times N_{b}$) time for 100 sums (in ms) performance penalty(in %)
SPA $(512\times 128)$ $6.456(8)$ 00
SPS $(512\times 128)$ $6.469(11)$ $-0.198538$
V100 TPRC $(512\times 128)$ $6.491(6)$ $-0.528162$
CU (unknown) $6.87697$ $-6.51331$
AO (fixed parameters) $872.004(22)$ $-28781.3$
SPA $(512\times 512)$ $3.019(22)$ 00
CU (unknown) $3.1553$ $-4.50533$
GH200 SPS $(512\times 512)$ $3.254(14)$ $-7.79916$
AO (fixed parameters) $738.687(37)$ $-24365.7$
TPRC $(512\times 256)$ $6.275(12)$ 00
CU (unknown) $6.378$ $-1.63577$
Mi250X SPA $(512\times 256)$ $6.394(32)$, $-1.90101$
SPS $(256\times 512)$ $6.552(23)$ $-4.4171$
AO (fixed parameters) $505.453(89)$ $-16640.9$
III-D Other approaches: Parallel sums with OpenMP

### v3

@@ Abstract
Run to run variability in parallel programs caused by floating-point non-associativity has been known to significantly affect reproducibility in iterative algorithms, due to accumulating errors. Non-reproducibility can critically affect the efficiency and effectiveness of correctness testing for stochastic programs. Recently, the sensitivity of deep learning training and inference pipelines to floating-point non-associativity has been found to sometimes be extreme. It can prevent certification for commercial applications, accurate assessment of robustness and sensitivity, and bug detection. New approaches in scientific computing applications have coupled deep learning models with high-performance computing, leading to an aggravation of debugging and testing challenges. Here we perform an investigation of the statistical properties of floating-point non-associativity within modern parallel programming models, and analyze performance and productivity impacts of replacing atomic operations with deterministic alternatives on GPUs. We examine the recently-added deterministic options in PyTorch within the context of GPU deployment for deep learning, uncovering and quantifying the impacts of input parameters triggering run to run variability and reporting on the reliability and completeness of the documentation. Finally, we evaluate the strategy of exploiting automatic determinism that could be provided by deterministic hardware, using the Groq $\text{LPU}^{\text{TM}}$ accelerator for inference portions of the deep learning pipeline. We demonstrate the benefits that a hardware-based strategy can provide within reproducibility and correctness efforts.
@@ I Introduction
Run to run variability of a program, despite identical inputs and software stack, is usually assumed to be negligible, even in heterogeneous parallel programming [1]. However, variability caused by floating-point non-associativity (FPNA) coupled with asynchronous parallel operations such as reductions can be substantial [2], especially for massively parallel programs using iterative stochastic routines, such as those implementing optimization algorithms like conjugate gradient [3]. It can also mask errors within threshold-based correctness testing schemes [4], which are often used in scientific computing programs such as molecular simulation [5, 6, 7], making debugging difficult. Deep neural network (DNN) training also involves iterative, stochastic algorithms coupled with non-linear activation functions. This combination has been found to cause extreme sensitivity to bit-level numerical changes, such as those caused by FPNA [8, 9, 10]. DNN inference can also suffer from such sensitivity, due to the influence of non-linear activation functions, although the effects may be reduced due to the absence of the compounding errors caused by the iterative training schemes. Recently, a number of studies have shown that this run to run variability in the full DNN training and inference pipeline can lead to unacceptably large differences in the predictions produced by a model. They have thwarted efforts to release reproducible commercial applications in safety-critical sectors such as medical diagnostics [11] and autonomous driving [8]. In addition, recent reports have found that all of the major deep learning (DL) software frameworks such as TensorFlow and PyTorch contain hundreds of bugs originating from all software levels, including environment misuse, incorrect assignment, and concurrency issues such as data races [12]; bugs in these frameworks can be disastrous [13], especially when they are silent [14], and high runtime variability can make it extremely difficult to detect bugs in these deep, multi-language, multi-level parallel stacks.
@@
Within scientific high-performance computing (HPC), the incorporation of DL into traditional approaches for simulation has become increasingly popular [15, 16, 17]. Molecular simulation, for example, has used DNN models for interatomic potentials, which promise quantum mechanical accuracy at the cost of simpler, empirical models; this translates to speedups of several orders of magnitude and the approach received the ACM Gordon Bell Award in 2020 [18]. Our preliminary tests using identical inputs for training and inference pipelines for these types of models revealed a level of non-reproducibility for prediction of forces that would be unacceptable for traditional quantum mechanical HPC programs [19]. Besides non-deterministic kernels within the deep learning framework itself, any external kernels programmed by the scientific simulation developers which contain non-determinism, that feed into training pipelines [18], can also introduce large runtime non-reproducibility.
@@
II Metrics for measuring the variability of non-deterministic functions
We defined metrics $V$ for quantifying the run to run variability in output values for implementations of functions with scalar and multidimensional (array) inputs and outputs, following a similar approach used in error analysis [20]. These are explained below. In all cases, $V=0$ if two implementations are bitwise identical, nonzero otherwise, and increasing as variability increases.
II-1 Scalar-valued outputs
We use $V_{s}(f)=1-\left|f_{\textsc{nd}}/f_{\textsc{d}}\right|$ to quantify the bitwise non-determinism between the outputs of two implementations of some function $f$, where the subscripts nd and d label the non-deterministic and deterministic implementations respectively.
II-2 Array outputs
To quantify the bitwise variability of a non-deterministic implementation of a function producing a multidimensional array output, we define two different metrics. Let two implementations of a function $f$ be $f_{1}$ and $f_{2}$, which produce as outputs the arrays $\mathbf{A}$ and $\mathbf{B}$, respectively, each with dimensions $d_{1},d_{2},\ldots,d_{k}$ and $D$ total elements. The first metric is the elementwise relative mean absolute variation (ERMV),
@@ II-2 Array outputs
The second metric, the count (C) variability, measures how many elements in the multidimensional array are different,
@@
III Programming deterministic parallel sums
In this section we first introduce FPNA impacts for the parallel sum, then demonstrate several deterministic programming solutions, for GPUs using CUDA, and also using OpenMP for CPU or GPU. We then examine the performance impacts of these solutions, as well as programming productivity considerations. Parallel sums, as we will show in the subsequent sections, contribute some of the most substantial sources of variability to PyTorch functions on GPUs, and are thus responsible for concerning non-determinism in real-world DL applications such as graph neural network (GNN) s, which make heavy use of them. A concerning connection to molecular physics and chemistry applications in scientific computing is the increased popularity of GNN s for surrogate models in these applications [21, 22], especially considering the stringent precision requirements for these applications, as discussed below.
Consider the sum $S_{\textsc{d}}=\sum_{i}^{n}x_{i}$ that we evaluate with a deterministic algorithm (the summation order is always the same). The simplest implementation adds the floating-point numbers (FP) $x_{i}$ in serial in the order they are stored. When parallelized with asynchronous operations and executed in an unspecified order, this is equivalent to applying a random permutation $P$ to the series $x_{i}$ before computing the serial sum, $S_{\textsc{nd}}=\sum_{i}^{n}x_{P(i)}$.
@@ III Programming deterministic parallel sums
To demonstrate the magnitude of the variability, we can generate lists $x_{i}$ of double precision floating point numbers (FP64) numbers of various lengths using Python, with, for example, $x_{i}\in\mathcal{N}(0,1)$ the normal distribution of zero mean and $\sigma=1$, and computed $S_{\textsc{d}}$ and $S_{\textsc{nd}}$ before and after applying a random permutation $P$ to $x_{i}$, repeating ten times; Table I shows the results. The variability can be larger than the tolerance of some of the correctness tests included in computational physics and chemistry programs; this same result obtains with a Boltzmann (exponential) distribution as well, which is the expected distribution for such calculations. For example, the quantum mechanics simulation program CP2K[5] uses a tolerance-based correctness testing scheme, and some of the tests have tolerances as low as $10^{-14}$ for values such as energy. This illustrates the problems that FPNA can introduce in correctness testing schemes for computational tasks with stringent accuracy requirements. In addition, such programs often employ iterative solvers such as conjugate gradient, leading to accumulating FP errors which can approach or exceed 20% after six or seven iterations using double precision, as has been shown for HPC systems such as massively multithreaded machines like the Cray XMT [3].
III-A Examples of deterministic parallel sum implementations on GPUs with CUDA/HIP
A difficulty for programming on GPUs is the absence of a global synchronization barrier at the kernel level. To avoid races one can use (i) a single atomicAdd function, (ii) the implicit global synchronization happening when multiple kernels are added to the same stream (iii) one GPU thread to calculate the sum, (iv) the __threadfence instruction.
1
2// implementation (AO) of the sum
3__global__ void reduce_atomic_only(double *sdata, int size, double *res) {
@@ III-A Examples of deterministic parallel sum implementations on GPUs with CUDA/HIP
7}
8
9__device__ int retirementCount = 0;
10// Implementation of the SPRG method
11__global__ void reduce_sprg(double *sdata, int size, double *res) {
12 int tid = threadIdx.x + blockIdx.x * blodkDim.x;
13 extern __shared__ double smem[];
14
@@
39}
Listing 1: Implementation of parallel sum using the atomicAdd
instruction, and deterministic implementation using pairwise algorithm.
Option (iii) is deterministic, but slow, as only one GPU thread computes the reduction. A code snippet can be found in the GitHub repository associated with this paper. Option (i) is only available for AMD GPUs (with HIP) in an unsafe compiler mode and is not considered for AMD GPUs here. This atomicAdd-only (AO) method is shown in Listing 1 for the CUDA version. This approach is the easiest to program (only a few lines of code), but is actually sequential, and yet the instruction order is runtime dependent, making it non-deterministic. Option (ii) and (iv) use a pairwise algorithm [20, 23] coupled to data blocking for distributing the data over the different thread blocks and specifying the order for summation. Here, each thread of a given block adds the elements $x_{i}$ in pairs $t_{i}=x_{i}+x_{i+n/2}$, $i=1,\cdots,n/2$, done in parallel on GPU. This process is repeated $log_{2}(n)$ times on $t_{i}$. We use the __syncthreads instruction after each step of the partial reduction for thread synchronization within the thread block. Shared memory is also used to improve data locality and performance. Partial results are then stored back in global memory. We tested three different methods for accumulating these partial results. The simple-pass-with atomicAdd (SPA), uses an atomicAdd instruction to accumulate all partial sums instead of storing the results back in memory. This is again a simple solution from the programmatic perspective, but the implementation is not deterministic. The two-passes-with-final-reduction-on-CPU (TPRC) method uses the property that two kernels launched on the same stream will execute sequentially following the submission order, introducing a synchronization barrier between them, or between a kernel and a data transfer between GPU and CPU. We choose the data transfer, and compute the final sum on CPU using the sequential recursive method. TPRC is deterministic, but more sensitive to compiler optimizations because of vectorization. The last approach uses an integer counter to keep track of the completed thread blocks. The sum function from the CUB/HIPCUB library (CU) uses a similar technique. Each thread block increments this variable atomically before exiting. The last block incrementing the counter is responsible for the remaining reduction, and can be performed in two ways. The single-pass-with-tree-reduction (SPTR) method uses the same block reduction algorithm than for the first stage while the single-pass-with-final-recursive-sum-on-GPU (SPRG) variant (see listing 1) uses the recursive sum. SPTR and SPRG are both deterministic by construction. SPRG and SPTR are only possible if we use the __threadfence instruction to avoid data race conditions. This instruction ensures that all writes issued by the calling thread are finished before the calling thread runs the next instruction, and notifies all the other running threads that the memory operation is over. The instruction is not a global synchronization barrier.
TABLE II: Different implementations of the parallel sum in CUDA. Non-deterministic implementations shown in red.
Method deterministic # of kernels synchronization methods
CU Yes - __threadfence
SPTR Yes 1 __threadfence
SPRG Yes 1 __threadfence
TPRC Yes 2 stream synchronization
SPA No 1 atomicAdd
AO No 1 atomicAdd
A summary of the main properties of each implementation is provided in Table II while codes snippets and full reference codes are provided in the GitHub repository [24].
III-B Other approaches: Parallel sums with OpenMP